OpenStudy (lilmike234):

What is the domain of the function shown below r(x)=x/x^2-9

OpenStudy (jdoe0001):

if you solve the denominator for 0, what do you get for "x"?

OpenStudy (usukidoll):

\[\large r(x) = \frac{x}{x^2-9}\] our function has a denominator so there is a restriction we can either factor \[\large x^2-9 \] due to difference of squares formula... or solve \[\large x^2-9=0\] for x. There should be 2 solutions.

OpenStudy (lilmike234):

(x:xnot equal to 3) ?

OpenStudy (usukidoll):

ummm we're missing another value for x hint \[\large x^2-9=0 \] difference of squares formula \[\large (x+3)(x-3)=0\] now solve for x

OpenStudy (lilmike234):

(x: all real numbers)?

OpenStudy (usukidoll):

solve x+3 = 0 and solve x-3=0

OpenStudy (lilmike234):

(x:xnot equal to -3) ?